# Trustworthy Transfer Learning: A Survey

### v1

@@ 4.3 Fairness
Fairness involves eliminating discrimination when training machine learning models (?, ?). In the legal domain, potential discrimination is defined as disparate treatment (triggered by intentionally treating an individual differently) and disparate impact (triggered by negatively affecting members of a protected group) (?). Motivated by this definition, different measures of algorithmic fairness have been proposed in machine learning communities, e.g., individual fairness (?), group fairness (?), etc. To be specific, individual fairness (?) maintains that similar individuals should be treated similarly. Group fairness (?, ?) ensures statistical parity among groups with sensitive attributes (e.g., race, gender, age). In the context of transfer learning, a fundamental concern is whether the fairness of a machine learning model can be transferred across domains under distribution shifts. Following (?, ?), the problem of fairness transfer can be formulated as follows.
Definition 9 (Fairness Transfer (?, ?))
@@
Given a source domain and a target domain, we denote the fairness violation measures as $\Delta^{*}_{S}(\cdot)$ for the source domain and $\Delta^{*}_{T}(\cdot)$ for the target domain. For any hypothesis $f\in\mathcal{F}$, algorithmic fairness can be transferred, if the following condition is satisfied:
$\Delta^{*}_{T}(f)\leq\Delta^{*}_{S}(f)+\delta$ (18)
@@
4.3.1 Group Fairness
Generally, group fairness (?, ?, ?) requires that different groups are treated equally. There are several commonly used group fairness metrics: demographic parity ? (?), equality of opportunity (?), and equalized odds (?). Following (?, ?), we formally define these metrics in a binary classification problem where $\mathcal{Y}=\{0,1\}$. Assuming there are two groups defined by binary sensitive attributes $A\in\{0,1\}$, fair machine learning seeks to ensure accurate predictions without bias against any particular group.
  • Demographic Parity (?): Demographic parity, also known as statistical parity, requires the same positive prediction ratio across groups with different sensitive attributes.
    $\mathrm{Pr}(\hat{Y}=1|A=0)=\mathrm{Pr}(\hat{Y}=1|A=1)$ (19)
    where $\hat{Y}$ denotes the random variable of the predicted class label and $\mathrm{Pr}(\cdot)$ represents the probability. This criterion implies that the decisions made by machine learning models should be independent of any sensitive attributes. However, it may be limited in scenarios where the base rates of the two groups differ, i.e., $\mathrm{Pr}({Y}=1|A=0)\neq\mathrm{Pr}({Y}=1|A=1)$ where $Y$ is the ground-truth class variable. In such cases, it is unrealistic to expect both model accuracy and demographic parity to be achieved simultaneously. Notably, ? (?) theoretically characterize the inherent trade-off between statistical parity and prediction accuracy, by providing a lower bound on group-wise prediction error for any fair predictor under demographic parity.
  • Equality of Opportunity (?): A machine learning model is considered fair under equality of opportunity if the false positive rates across groups are equal.
    $\mathrm{Pr}(\hat{Y}=1|A=0,Y=0)=\mathrm{Pr}(\hat{Y}=1|A=1,Y=0)$ (20)
@@ 4.3.1 Group Fairness
  • Equalized Odds (?): A machine learning model is considered fair under equalized odds if both the false positive rates and false negative rates across groups are equal.
    $\mathrm{Pr}(\hat{Y}=1|A=0,Y=0)$ $=\mathrm{Pr}(\hat{Y}=1|A=1,Y=0)$ (21)
@@
Fair transfer learning that integrates the aforementioned group fairness criteria has been studied in recent years (?, ?). For example, ? (?) and ? (?) propose learning fair intermediate representations by encoding the data as accurately as possible while obscuring information about sensitive attributes. They demonstrate the transferability of these fair representations across different tasks. Later, ? (?) empirically investigate the connections between compound distribution shifts (e.g., the co-occurrence of demographic, covariate, and label shifts) and fairness transfer in real-world medical applications via a joint causal framework (?). Furthermore, ? (?) provide a generic Lipshitz upper bound for group fairness when the underlying distribution shifts (e.g., covariate shift or label shift between source and target domains) are constrained. Specifically, most existing works (?, ?, ?) dive into understanding the transferability of fairness, by considering various learning scenarios based on the availability of class labels and sensitive attribute information in the target domain. Figure 11 illustrates three scenarios for group fairness transfer when training data are available in the target domain. Another related scenario is the domain generalization (?) where no target samples are available during training.
Figure 11: Transferability of group fairness across domains. (a) The target domain has labeled samples with sensitive attributes. (b) The target domain has unlabeled samples with sensitive attributes. (c) The target domain has only unlabeled samples without sensitive attributes.
@@
    Labeled target samples with sensitive attributes: Assuming that the target domain contains a few labeled samples with sensitive attributes, ? (?) provide the generalization error bounds of group fairness (e.g., equality of opportunity, and equalized odds) in the target domain in terms of fairness-aware distribution discrepancy between source and target domains. ? (?, ?) theoretically show the generalization error bound of group fairness (e.g., demographic parity) across domains from the perspective of multi-task learning via low-rank matrix factorization or parameter decoupling. Similarly, ? (?) propose a fair meta-learning algorithm to transfer the fairness across domains.
@@
    Unlabeled target samples with sensitive attributes: When the target domain has only unlabeled samples with sensitive attributes, ? (?) propose minimizing both the expected log-loss and the pseudo-label aware fairness penalty over the worst-case approximation of the target distribution to mitigate covariate shifts across domains and ensure fairness (e.g., demographic parity, equality of opportunity, and equalized odds) in the target domain. ? (?) leverage representation matching across sensitive groups to enforce fairness and sample reweighting to mitigate covariate shifts across domains. Inspired by the theory of self-training (?, ?), ? (?) theoretically analyze the transferability of group fairness across domains based on the consistency loss of a machine learning model under input transformations. Then they propose a self-training algorithm with fair consistency regularization to improve fairness transfer in the presence of subpopulation shifts. In contrast, ? (?) formalize the notion of correlation shift over labels and sensitive attributes and employ a weighted sampling strategy in data preprocessing to mitigate correlation shifts across domains.
@@
    Only unlabeled target samples with missing sensitive attributes: ? (?) study a more general learning scenario where the target domain is associated with only unlabeled samples with missing sensitive attributes. To improve group fairness (e.g., demographic parity) in the target domain, they develop fairness-guided sample reweighting approaches by enforcing the similarity of group-wise weighting scores across all pairs of groups.
@@
    No target samples: An extreme situation occurs when no target samples are available, commonly referred to as domain generalization or out-of-distribution generalization (?, ?). In this scenario, only source domain data is provided to learn a fair predictor for unseen target domains. To solve this problem, ? (?) develop a causal inference framework to minimize the worst-case prediction error under group fairness constraints. Similarly, ? (?) focus on optimizing a fair predictor by minimizing the worst-case error across weighted combinations of the training data. Later, ? (?) theoretically derive the upper bounds on generalization error and unfairness in the target domain in terms of source error/unfairness, the domain discrepancy among source domains, and the domain discrepancy between source and unseen target domains. Motivated by this theoretical analysis, they propose an invariant representation learning algorithm to improve the transfer of fairness and accuracy via density matching.
4.3.2 Individual Fairness
@@
Individual fairness requires that similar individuals (in the input space) should receive similar decision outcomes (in the output space) (?, ?). Individuals are similar if their only differences lie in protected attributes or features related to those attributes. Mathematically, ? (?) formalize this notion using $L$-Lipschitz continuity of a function $f:\mathcal{X}\to\mathcal{Y}$. For all $x_{1},x_{2}\in\mathcal{X}$, the following holds
$d_{\mathcal{Y}}\left(f\left(x_{1}\right),f\left(x_{2}\right)\right)\leq L\cdot d_{\mathcal{X}}\left(x_{1},x_{2}\right)$ (22)
@@
where $L$ is a constant. Here, $d_{\mathcal{X}}$ and $d_{\mathcal{Y}}$ represent the distance metrics in the input space and output space, respectively. Recently, ? (?) investigate the connections between individual fairness and knowledge transferability in unsupervised domain adaptation/generalization scenarios. They show that (i) enforcing individual fairness (e.g., graph Laplacian regularizer (?)) can theoretically improve the generalization performance of a learning function under the covariate shift assumption, and (ii) invariant representation learning commonly used in existing domain adaptation algorithms (?) can improve individual fairness. The follow-up work (?) introduces a two-stage pre-training and fine-tuning framework based on graph Laplacian regularize to enforce the transferability of individual fairness across domains. Besides, ? (?) propose an end-to-end framework to learn individually fair representations with provable certification and demonstrate the transferability of individual fairness using the learned representation. ? (?) further study the certification of distributional individual fairness (?), which enforces the individual fairness within a $\gamma$-Wassertein ball of the empirical distribution over a finite set of observed individuals. The proposed distributional individual fairness regularization explicitly enables the transferability of individual fairness under in-the-wild distribution shifts.
4.4 Transparency
@@
Transparency helps non-experts understand the decision-making process of a machine learning model and the confidence level of the model in making decisions (?). For example, interpretability and explainability have recently been studied to enhance transparency, by designing a simpler and more interpretable model (?, ?) or providing post-hoc explanations for existing black-box models (?). As a complementary metric of transparency, uncertainty quantification (?) illustrates the prediction confidence of a trained model. As a result, we study two major questions behind transparent transfer learning: what knowledge is being transferred in transfer learning, and how to quantify the uncertainty of transfer learning models.
4.4.1 Interpretability/Explanability
@@
Despite the promising performance of transfer learning techniques in a range of applications, there is still limited understanding of which data and model architecture components contribute to successful knowledge transfer across domains. To bridge this gap, ? (?) demonstrate that for neural networks pre-trained on ImageNet data set (?), modules in the lower layers are responsible for capturing general features (e.g., Gabor and color blob features in images), while the higher-layer modules tend to encode task-specific semantic features. ? (?) further support this finding from the perspective of module criticality (?). Besides, they also reveal that both feature-reuse and low-level data statistics are crucial for successful knowledge transfer. More recently, ? (?) establish the connections between fine-tuned neural layers and types of distribution shifts (shown in Figure 12). They find that fine-tuning the first block is most effective for input-level shifts (such as image corruption), intermediate blocks excel at feature-level shifts (like shifts in entity subgroups), and tuning the last layer is best for output-level shifts (such as spurious correlations between gender and hair color). ? (?) also investigate transfer learning for medical imaging. They show that using a larger pre-trained ImageNet model does not significantly improve performance compared to smaller lightweight convolutional networks. Additionally, it is observed that transfer learning provides feature-independent benefits, such as improved weight scaling and faster convergence. This is consistent with observations from (?, ?).
@@
Figure 12: Illustration of surgical fine-tuning (adapted from (?)), where the selected fine-tuning blocks are correlated with the types of distribution shifts between source and target domains.
@@
In addition to understanding the transferability of pre-trained models, recent efforts have been devoted to exploring the explanations of distribution shifts across domains in the distribution space. There are two major frameworks for distribution shift explanations, including interpretable transportation mapping (?, ?) and natural languages (?, ?, ?). Specifically, on one hand, ? (?, ?) explain distribution shifts using interpretable transportation maps indicating how the source distribution can move to the target distribution in the distribution space. The crucial idea is to leverage optimal transport to find the optimal transportation map from user-defined interpretable candidates. ? (?) further propose a group-aware shift explanation framework to rectify the group irregularities when explaining distribution shifts. On the other hand, ? (?) develop a GSCLIP system to explain distribution shifts of different image data sets in natural language. This system generates human-understandable natural language descriptions of distribution shifts as candidate explanations, and then quantitatively evaluates these candidates to identify the most reasonable ones. ? (?) study the explanations for text distribution shifts through natural languages. They prompt GPT-3 (?) to generate candidate explanations and then employ a verifier neural network to re-rank these explanations. Similarly, ? (?) leverage visual-language models to generate candidate difference descriptions from image sets and then re-rank these candidates based on their effectiveness in distinguishing the two sets. In contrast, based on graphical causal models, ? (?) propose a Shapley value (?) framework to quantify the attribution for each causal mechanism for distribution shifts. The follow-up work (?) further studies connections between model performance changes across domains and interpretable distribution shifts via Shapley values.
4.4.2 Uncertainty Quantification

### v2

@@ 4.3. Fairness
Fairness involves eliminating discrimination when training machine learning models (eshete2021making; castelnovo2022clarification). In the legal domain, potential discrimination is defined as disparate treatment (triggered by intentionally treating an individual differently) and disparate impact (triggered by negatively affecting members of a protected group) (pessach2023algorithmic). Motivated by this definition, different measures of algorithmic fairness have been proposed in machine learning communities, e.g., individual fairness (DworkHPRZ12), group fairness (FeldmanFMSV15), etc. To be specific, individual fairness (DworkHPRZ12) maintains that similar individuals should be treated similarly. Group fairness (FeldmanFMSV15; HardtPNS16) ensures statistical parity among groups with sensitive attributes (e.g., race, gender, age). In the context of transfer learning, a fundamental concern is whether the fairness of a machine learning model can be transferred across domains under distribution shifts. Following (schumann2019transfer; chen2022fairness), the problem of fairness transfer can be formulated as follows.
Definition 4.3 (Fairness Transfer (schumann2019transfer; chen2022fairness)).
@@
Given a source domain and a target domain, we denote the fairness violation measures as $\Delta^{*}_{S}(\cdot)$ for the source domain and $\Delta^{*}_{T}(\cdot)$ for the target domain. For any hypothesis $f\in\mathcal{F}$, algorithmic fairness can be transferred if the following condition is satisfied:
(18) $\Delta^{*}_{T}(f)\leq\Delta^{*}_{S}(f)+\delta$
@@
4.3.1. Group Fairness
Generally, group fairness (FeldmanFMSV15; HardtPNS16; castelnovo2022clarification) requires that different groups are treated equally. There are several commonly used group fairness metrics: demographic parity (FeldmanFMSV15), equality of opportunity (HardtPNS16), and equalized odds (HardtPNS16). Following (MadrasCPZ18; schumann2019transfer), we formally define these metrics in a binary classification problem where $\mathcal{Y}=\{0,1\}$. Assuming there are two groups defined by binary sensitive attributes $A\in\{0,1\}$, fair machine learning seeks to ensure accurate predictions without bias against any particular group.
  • Demographic Parity (FeldmanFMSV15): Demographic parity, also known as statistical parity, requires the same positive prediction ratio across groups with different sensitive attributes.
    (19) $\mathrm{Pr}(\hat{Y}=1|A=0)=\mathrm{Pr}(\hat{Y}=1|A=1)$
    where $\hat{Y}$ denotes the random variable of the predicted class label and $\mathrm{Pr}(\cdot)$ represents the probability. This criterion implies that the decisions made by machine learning models should be independent of any sensitive attributes. However, it may be limited in scenarios where the base rates of the two groups differ, i.e., $\mathrm{Pr}({Y}=1|A=0)\neq\mathrm{Pr}({Y}=1|A=1)$ where $Y$ is the ground-truth class variable. In such cases, it is unrealistic to expect both model accuracy and demographic parity to be achieved simultaneously. Notably, (ZhaoG19) theoretically characterize the inherent trade-off between statistical parity and prediction accuracy, by providing a lower bound on group-wise prediction error for any fair predictor under demographic parity.
  • Equality of Opportunity (HardtPNS16): A machine learning model is considered fair under equality of opportunity if the false positive rates across groups are equal.
    (20) $\mathrm{Pr}(\hat{Y}=1|A=0,Y=0)=\mathrm{Pr}(\hat{Y}=1|A=1,Y=0)$
@@ 4.3.1. Group Fairness
  • Equalized Odds (HardtPNS16): A machine learning model is considered fair under equalized odds if both the false positive rates and false negative rates across groups are equal.
    (21) $\mathrm{Pr}(\hat{Y}=1|A=0,Y=0)$ $=\mathrm{Pr}(\hat{Y}=1|A=1,Y=0)$
@@
Fair transfer learning that integrates the aforementioned group fairness criteria has been studied in recent years (giguere2022fairness; dutt2024fairtune). For example, (zemel2013learning; MadrasCPZ18) propose learning fair intermediate representations by encoding the data as accurately as possible while obscuring information about sensitive attributes. They demonstrate the transferability of these fair representations across different tasks. Later, (SchrouffHKASOBR22) empirically investigate the connections between compound distribution shifts (e.g., the co-occurrence of demographic, covariate, and label shifts) and fairness transfer in real-world medical applications via a joint causal framework (mooij2020joint). Furthermore, (chen2022fairness) provide a generic Lipshitz upper bound for group fairness when the underlying distribution shifts (e.g., covariate shift or label shift between source and target domains) are constrained. Specifically, most existing works (coston2019fair; biswas2021ensuring; zhao2024algorithmic) dive into understanding the transferability of fairness, by considering various learning scenarios based on the availability of class labels and sensitive attribute information in the target domain. Fig. 11 illustrates three scenarios for group fairness transfer when training data are available in the target domain. Another related scenario is the domain generalization (pham2023fairness) where no target samples are available during training.
(a) Fairness transfer with $(X_{T},A_{T},Y_{T})$
(b) Fairness transfer with $(X_{T},A_{T})$
(c) Fairness transfer with $X_{T}$
Figure 11. Transferability of group fairness across domains. (a) The target domain has labeled samples with sensitive attributes. (b) The target domain has unlabeled samples with sensitive attributes. (c) The target domain has only unlabeled samples without sensitive attributes.
@@
    Labeled target samples with sensitive attributes: Assuming that the target domain contains a few labeled samples with sensitive attributes, (schumann2019transfer) provide the generalization error bounds of group fairness (e.g., equality of opportunity, and equalized odds) in the target domain in terms of fairness-aware distribution discrepancy between source and target domains. (oneto2020learning; oneto2020exploiting) theoretically show the generalization error bound of group fairness (e.g., demographic parity) across domains from the perspective of multi-task learning via low-rank matrix factorization or parameter decoupling. Similarly, (slack2020fairness) propose a fair meta-learning algorithm to transfer the fairness across domains.
@@
    Unlabeled target samples with sensitive attributes: When the target domain has only unlabeled samples with sensitive attributes, (rezaei2021robust) propose minimizing both the expected log-loss and the pseudo-label aware fairness penalty over the worst-case approximation of the target distribution to mitigate covariate shifts across domains and ensure fairness (e.g., demographic parity, equality of opportunity, and equalized odds) in the target domain. (havaldar2024fairness) leverage representation matching across sensitive groups to enforce fairness and sample reweighting to mitigate covariate shifts across domains. Inspired by the theory of self-training (wei2021theoretical; cai2021theory), (AnCDH22) theoretically analyze the transferability of group fairness across domains based on the consistency loss of a machine learning model under input transformations. Then they propose a self-training algorithm with fair consistency regularization to improve fairness transfer in the presence of subpopulation shifts. In contrast, (roh2023improving) formalize the notion of correlation shift over labels and sensitive attributes and employs a weighted sampling strategy in data preprocessing to mitigate correlation shifts across domains.
@@
    Only unlabeled target samples with missing sensitive attributes: (coston2019fair) study a more general learning scenario where the target domain is associated with only unlabeled samples with missing sensitive attributes. To improve group fairness (e.g., demographic parity) in the target domain, they develop fairness-guided sample reweighting approaches by enforcing the similarity of group-wise weighting scores across all pairs of groups.
@@
    No target samples: An extreme situation occurs when no target samples are available, commonly referred to as domain generalization or out-of-distribution generalization (BlanchardLS11; gulrajani2021in). In this scenario, only source domain data is provided to learn a fair predictor for unseen target domains. To solve this problem, (singh2021fairness) develop a causal inference framework to minimize the worst-case prediction error under group fairness constraints. Similarly, (mandal2020ensuring) focus on optimizing a fair predictor by minimizing the worst-case error across weighted combinations of the training data. Later, (pham2023fairness) derive the theoretical upper bounds on generalization error and unfairness in the target domain in terms of source error/unfairness, the domain discrepancy among source domains, and the domain discrepancy between source and unseen target domains. Motivated by this theoretical analysis, they propose an invariant representation learning algorithm to improve the transfer of fairness and accuracy via density matching.
4.3.2. Individual Fairness
@@
Individual fairness requires that similar individuals (in the input space) should receive similar decision outcomes (in the output space) (DworkHPRZ12; zemel2013learning). Individuals are similar if their only differences lie in protected attributes or features related to those attributes. Mathematically, (DworkHPRZ12) formalize this notion using $L$-Lipschitz continuity of a function $f:\mathcal{X}\to\mathcal{Y}$. For all $x_{1},x_{2}\in\mathcal{X}$, the following holds
(22) $d_{\mathcal{Y}}\left(f\left(x_{1}\right),f\left(x_{2}\right)\right)\leq L\cdot d_{\mathcal{X}}\left(x_{1},x_{2}\right)$
@@
where $L$ is a constant. Here, $d_{\mathcal{X}}$ and $d_{\mathcal{Y}}$ represent the distance metrics in the input space and output space, respectively. Recently, (MukherjeePYS22) investigate the connections between individual fairness and knowledge transferability in unsupervised domain adaptation/generalization scenarios. They show that (i) enforcing individual fairness (e.g., graph Laplacian regularizer (KangHMT20)) can theoretically improve the generalization performance of a learning function under the covariate shift assumption, and (ii) invariant representation learning commonly used in existing domain adaptation algorithms (DANN) can improve individual fairness. Besides, (ruoss2020learning) propose an end-to-end framework to learn individually fair representations with provable certification and demonstrates the transferability of individual fairness using the learned representation. (WickerPW23) further study the certification of distributional individual fairness (Yurochkin2020Training), which enforces the individual fairness within a $\gamma$-Wassertein ball of the empirical distribution over a finite set of observed individuals. The proposed distributional individual fairness regularization explicitly enables the transferability of individual fairness under in-the-wild distribution shifts.
4.4. Transparency
@@
Transparency helps non-experts understand the decision-making process of a machine learning model and the confidence level of the model in making decisions (Varshney2022). For example, interpretability and explainability have recently been studied to enhance transparency, by designing a simpler and more interpretable model (KohL17; Ribeiro0G16) or providing post-hoc explanations for existing black-box models (SelvarajuCDVPB17). As a complementary metric of transparency, uncertainty quantification (BhattAZLSFMKSTN21) illustrates the prediction confidence of a trained model. As a result, we study two major questions behind transparent transfer learning: what knowledge is being transferred in transfer learning, and how to quantify the uncertainty of transfer learning models.
4.4.1. Interpretability/Explanability
@@
Despite the promising performance of transfer learning techniques in a range of applications, limited effort has been devoted to understanding which data and model architecture components contribute to successful knowledge transfer across domains. To bridge this gap, (YosinskiCBL14) demonstrate that for neural networks pre-trained on ImageNet data set (deng2009imagenet), modules in the lower layers are responsible for capturing general features (e.g., Gabor and color blob features in images), while the higher-layer modules tend to encode task-specific semantic features. (NeyshaburSZ20) further support this finding from the perspective of module criticality (Chatterji2020The). Besides, they also reveal that both feature reuse and low-level data statistics are crucial for successful knowledge transfer. More recently, (lee2023surgical) establish the connections between fine-tuned neural layers and types of distribution shifts (shown in Fig. 12). They find that fine-tuning the first block is most effective for input-level shifts (such as image corruption), intermediate blocks excel at feature-level shifts (like shifts in entity subgroups), and tuning the last layer is best for output-level shifts (such as spurious correlations between gender and hair color). (raghu2019transfusion) also investigate transfer learning for medical imaging. They show that using a larger pre-trained ImageNet model does not significantly improve performance compared to smaller lightweight convolutional networks. Additionally, it is observed that transfer learning provides feature-independent benefits, such as improved weight scaling and faster convergence. This is consistent with observations from (kornblith2019better; he2019rethinking).
@@
Figure 12. Illustration of surgical fine-tuning (adapted from (lee2023surgical)), where the selected fine-tuning blocks are correlated with the types of distribution shifts between source and target domains.
@@
In addition to understanding the transferability of pre-trained models, recent efforts have been devoted to exploring the explanations of distribution shifts across domains in the distribution space. There are two major frameworks for distribution shift explanations, including interpretable transportation mapping (KulinskiI23; stein2023rectifying) and natural languages (dunlap2024describing; zhu2022gsclip; zhong2023goal). Specifically, on one hand, (KulinskiI23; kulinski2022towards) explain distribution shifts using interpretable transportation maps indicating how the source distribution can move to the target distribution in the distribution space. The crucial idea is to leverage optimal transport to find the optimal transportation map from user-defined interpretable candidates. (stein2023rectifying) further propose a group-aware shift explanation framework to rectify the group irregularities when explaining distribution shifts. On the other hand, (zhu2022gsclip) develop a GSCLIP system to explain distribution shifts of different image data sets in natural language. This system generates human-understandable natural language descriptions of distribution shifts as candidate explanations, and then quantitatively evaluates these candidates to identify the most reasonable ones. (zhong2022describing) study the explanations for text distribution shifts through natural languages. They prompt GPT-3 (GPT3) to generate candidate explanations and then employ a verifier neural network to re-rank these explanations. Similarly, (dunlap2024describing) leverage visual-language models to generate candidate difference descriptions from image sets and then re-rank these candidates based on their effectiveness in distinguishing the two sets. In contrast, based on graphical causal models, (budhathoki2021did) propose a Shapley value framework to quantify the attribution for each causal mechanism for distribution shifts. The follow-up work (zhang2023did) further explores connections between model performance changes across domains and interpretable distribution shifts via Shapley values.
4.4.2. Uncertainty Quantification